40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 51     Entry time: Mon Jul 18 18:05:46 2016
Author: Gabriele 
Type: Optics 
Category: Daily Progress 
Subject: Optical setup 

Small modifications to the optical setup:

  • added a 1m focal lenght lens (AR coated for 1064nm...) right before the beam enters the viewport, to reduce the beam spot size on the QPD
  • added a 0.3 ND right after the laser, to avoid saturation of the QPD (about 75 microwatts total on the QPD)

ELOG V3.1.3-